OpenStudy (anonymous):

find the perfect square of x^2-4-2=0

OpenStudy (anonymous):

x^2-4x-2=0

OpenStudy (anonymous):

The pattern you are trying to match is\[x^2-2ax+a^2=(x-a)^2\]

OpenStudy (anonymous):

How do i start the problem out?

OpenStudy (anonymous):

Look at the pattern I gave you , and the pattern of the problem. Can you find a value of a to plug into my pattern that will give your equation?